I need some help.
If a homemade rocket is propelled upward from ground level with an initial velocity of 112 feet per second, after t seconds, its height h in feet is given by the formula h=-16t^2+112t. After how many seconds will its height be 160 feet?

Answer by ewatrrr(24785)

2 sec
160=-16t^2+112t
t^2 -7t + 10 = 0
(t -2)(t-7) = 0, t = 2sec it will first be at 160ft